# SonicID: User Identification on Smart Glasses with Acoustic Sensing

### v1

@@ 6.2. Impact of Different Amounts of Training Data
Figure 6. Average TPR, FPR and BAC across 24 Participants with Different Amounts of Training Data.
In the evaluation results above, we adopted 4 minutes of training data from each participant, which leads to satisfactory performance. However, in practice, the fewer data a new user needs to provide to register on the device, the more user-friendly the system will be. Therefore, we conducted another experiment to explore the impact of different amounts of training data on the system performance. The average TPR, FPR and BAC among 24 participants with different amounts of training data is illustrated in Fig. 6. As shown in the figure, both TPR and FPR decreased when we reduced the amount of data to train the model. From 4 minutes of data to 1 minute of data, TPR drops from 96.5% to 90.3% on Day 1, from 91.3% to 72.3% on Day 2 and from 80.7% to 50.8% on Day 3 respectively. In the meantime, the FPR also decreases from 4.1% to 1.6% on Day 1, from 4.5% to 2.3% on Day 2 and from 4.2% to 1.9% on Day 3. Hence, BAC changes from 96.2% to 94.3% on Day 1, from 93.4% to 85% on Day 2 and from 88.3% to 74.4% on Day 3 correspondingly. These results indicates that the real users have more difficulty authenticating themselves into the device but the system protects the device from unknown users better, with fewer amounts of training data. Since FPR is a more important metric in authentication systems as discussed above, we can just collect 1 minute of training data when a new user registers on the device, leading to a TPR of 90.3% and a FPR of 1.6%. With time passing by, the authentication mechanism might tend to fail more frequently when the registered users want to use the device. However, new training data can be continuously collected every time when the user wears the device and makes some operations. The model will be updated accordingly to make sure that the registered users have no problem getting access to the device.
6.3. Comparison of Two Sides of Sensors
When we designed the prototype for the SonicID system, we instrumented one pair of speaker and microphone on each side of the glasses to collect complete biometric information from users’ face. However, many commercial smart glasses only have the speaker and microphone on one side of the device. As a result, we would like to experiment on how the system performs when only one side of sensors are utilized. As designed in Sec. 3, four speaker-microphone channels are input into the deep learning model when two pairs of speakers and microphones are deployed. If we only want to use on pair of speaker and microphone on either right side or left side of the glasses, we can just input one speaker-microphone channel related to this pair of speaker and microphone into the model. The experiment results are demonstrated in Tab. 2.
@@ 6.3. Comparison of Two Sides of Sensors
Table 2. Average TPR, FPR and BAC across 24 Participants with Different Channels of Signals. Data Format: TPR/FPR/BAC.
@@
Both Sides 96.5%/4.1%/96.2% 91.3%/4.5%/93.4% 80.7%/4.2%/88.3%
Right Side 91.8%/3.5%/94.2% 69.7%/4.1%/82.8% 61.1%/3.9%/78.6%
Left Side 81.1%/2.5%/89.3% 38.7%/3.1%/67.8% 21.8%/2.4%/59.7%
@@ 6.4. Usability
A questionnaire was distributed to participants after the user study to collect their demographic information and feedback on the prototyped wearable device. First, the participants evaluated the overall experience of wearing the device by answering the question "How comfortable is this wearable device to wear around the face? (0 most uncomfortable, 5 most comfortable)". Among 24 participants, the average rating for this question is 3.6 (std=0.7), indicating that the device with the SonicID system is overall comfortable to wear around the face. Next we specifically asked participants to evaluate the weight of the device with the question "How acceptable do you find the weight of our wearable device? (0 most unacceptable, 5 most acceptable)". The average rating is 3.8 (std=0.9), validating that the weight of the device does not cause much burden on users.
In the SonicID system, we utilized two ultrasonic signals in the frequency ranges over 18 kHz. Theoretically, most people are not capable of hearing the sounds emitted from the system. However, due to the limitation of the hardware components, there might be some frequency leakage into the lower frequency range and some users might be able to hear the sound. As a result, we asked the participants to also answer the question "Can you hear the sound emitted from our system?". 18 out of 24 participants answered "No" to this question while the other 6 participants reported "Yes". For those participants who answered "Yes" to this question, a follow-up question "If yes, how comfortable do you feel about the emitted sound? (0 most uncomfortable, 5 most comfortable)" was asked. The average rating for this question among 6 participants is 4.3 (std=0.5). Even though some users might hear the sound from the system, they generally find this sound not bothering them a lot. Since the echo profile we use as an instance to authenticate users is as short as 0.12 s as discussed in Sec. 3.4, we do not expect the sound to trouble the users continuously. Furthermore, it is possible to lower the strength of the transmitted signals to make them less noticeable by users.
@@ 7.1. Power Consumption
We wanted to evaluate the power consumption of the SonicID system in operation. For this purpose, we measured the current that flowed through system with a current ranger [^5] https://lowpowerlab.com/guide/currentranger/ when the system was operating. The current was measured at 165.4 mA with a voltage at 3.3 V. This gave us an average power consumption of 545.8 mW. Considering that SonicID only needs an instance of 0.12 s to authenticate users every time, it consumes 19.8 mAs of energy for each authentication trial. If SonicID is deployed on Meta Ray-Ban with a battery capacity of 154 mAh, the authentication system can be activated around 28,000 times in theory, if SonicID is used alone.
@@ 7.4. Potential Methods to Improve Performance
As evaluated in Sec. 6, although the FPR is consistently below 5% across several days in evaluation, the TPR reduces on Day 2 and Day 3 compared with Day 1 for some participants. We believe this is due to several reasons. Firstly, the positive samples we collected from each participant every day were just from the 5 minutes of data. The small size of the data can lead to fluctuations in the evaluation results in TPR. Secondly, we noticed that TPR is especially low for P20 who has long hair and bangs. The hair can sometimes cover the system and prevent the participant from putting the glasses back to the same position on their head every time they remount the device. In future, we plan to explore several potential methods to improve the performance of the system.
@@ 7.4.1. Explore Movement-based Authentication
In Sec. 6, we removed information of blinking and other movements from the data during training and testing to guarantee a valid facial scanning outcome. However, this information can also be taken advantage of to authenticate users. For instance, the blinking information alone can be exploited as a unique biometric pattern because different users blink differently. Furthermore, we can explore other movement-based authentication methods to add another layer of protection. Users of our system can select a specific facial gesture (e.g. smile, tongue out, blink, etc.) as their password to the system. Our system recognizes both their password and their biometric information when performing this password to authenticate users. This kind of two-layer authentication systems make the system more secure and are possible to implement given that there are already some work validating that deploying acoustic sensing on wearables can help track users’ facial movements (Li et al., 2022; Zhang et al., 2023; Xie et al., 2021; Appear], 2024). SonicID can potentially be integrated into these existing systems to realize user authentication without adding extra hardware.

### v3

@@ 6.2. Impact of Different Amounts of Training Data
Figure 5. Average TPR, FPR and BAC across 40 Participants with Different Amounts of Training Data.
In the evaluation results above, we adopted 4 minutes of training data from each participant, which leads to satisfactory performance. However, in practice, the fewer data a new user needs to provide to register on the device, the more user-friendly the system will be. Therefore, we conducted another experiment to explore the impact of different amounts of training data on the system performance. To evaluate this factor, we discarded the last 2.5 seconds, 5 seconds and 7.5 seconds of data from each 10-second session and still picked ten most static 0.06-second instances from the remaining data in each session to train the model. In practice, this reduces the amount of training data that each new user needs to provide from 4 minutes to 3 minutes, 2 minutes and 1 minute respectively. The average TPR, FPR and BAC among 40 participants with different amounts of training data is illustrated in Fig. 5. As shown in the figure, all of the three metrics remain similar across three days when the amount of training data decreases from 4 minutes to 1 minute. We ran three one-way repeated measures ANOVA tests among TPR with different amounts of training data for Day 1, Day 2 and Day 3 and did not find a statistically significant difference on any of these days ($F(3,156)=1.37,p=0.26>0.05$ for Day 1, $F(3,156)=0.38,p=0.77>0.05$ for Day 2, and $F(3,156)=1.82,p=0.15>0.05$ for Day 3). The same ANOVA tests were performed among FPR and no statistically significant difference was found on any of the three days ($F(3,156)=1.32,p=0.27>0.05$ for Day 1, $F(3,156)=2.35,p=0.08>0.05$ for Day 2, and $F(3,156)=1.33,p=0.27>0.05$ for Day 3). The statistical analysis validates that the system performance maintains satisfactory with the amount of training data as little as 1 minute from each new user. Therefore, we adopted 1 minutes of data from each participant on each day to train the model and evaluate the performance in all the experiments below.
6.3. Leave-One-Day-Out Evaluation
In the experiments above, we utilized participants’ data on Day 1 to train the model and the data on Day 2 and Day 3 to test the model. In order to eliminate the impact of random factors among different days, we conducted leave-one-day-out experiments by using data from Day 2 and Day 3 to train the model respectively while the data from the other two days was used for evaluation. The experiment results are demonstrated in Tab. 2.
@@ 6.3. Leave-One-Day-Out Evaluation
Table 2. Average TPR, FPR and BAC across 40 Participants with Models Trained on 1-Minute Data from Different Days. Data Format: TPR/FPR/BAC. Results with Training and Testing Data from the Same Day are in Bold.
Training Data Source Day 1 Day 2 Day 3
Train on Day 1 97.4%/4.3%/96.6% 80.8%/4.1%/88.4% 75.2%/3.8%/85.7%
Train on Day 2 80.0%/3.7%/88.1% 96.5%/3.6%/96.5% 82.7%/3.6%/89.6%
Train on Day 3 76.4%/3.8%/86.3% 82.7%/3.2%/89.7% 98.2%/3.1%/97.5%
By averaging the leave-one-day-out evaluation results, the obtained average TPR, FPR and BAC when the system is evaluated on the same day as the training data is collected are 97.4%, 3.7% and 96.9% while those metrics are 79.6%, 3.7% and 88.0% on a different day. As a result, SonicID achieves a balanced accuracy over 95% across different remounting sessions and that over 85% across different days with a false positive rate under 4%. These results again validated the capability and stability of SonicID especially in terms of protecting users’ personal devices from attackers.
6.4. Performance Boost across Days
We noticed that SonicID achieves promising performance at protecting the device from bad actors, even across several days. However, the system starts to experience more failure of authenticating the real users into the device across days, especially for several specific participants. This can be attributed to the lack of training samples from real users as we only trained the model with 1 minutes of data collected on the same day. However, the scanned acoustic features of users’ faces can vary across days depending on their facial conditions such as how moist their faces are, the usage of makeups, the length of their beard, the coverage of their hair and so on. Usually the variation of the these factors should be subtle. However, under certain circumstances, these factors might significantly affect the reflection of the acoustic signals, leading to an unsatisfactory success rate of authentication. To validate our assumption, we plotted the reflected signals in the frequency domain from P25’s data and P30’s data on all three days. These two participants are selected because P30 has the lowest TPR on Day 3 and a low TPR on Day 2 (37.7%) while P25 has a TPR of 100% across all three days, if the model is trained on data from Day 1. To compare the change of signal reflection across days more clearly, we extracted the envelope of these signals to remove unnecessary details. The envelope signals are demonstrated in Fig. 6.
(a) Reflected Signals from P25
(b) Reflected Signals from P30
Figure 6. Envelope of Reflected Signals in Frequency Domain from P25’s Data and P30’s Data across Three Days.
As shown in the figure, there are subtle changes in the reflected signals across three days for P25 but overall they remain similar to each other. However, for P30, the signals largely change across different days. Specifically, signal reflection on Day 1 is quite different from the other two days. This explains why TPR of SonicID across days for P30 is worse than P25. As stated in Sec. 6.1, we believe that this may be owing to the long hair of P30 restraining the device from being put back to the same position every time of remounting. To address this issue, we decided to further fine-tune the individual model for participants with trouble logging into their device on a different day. The fine-tuning process of SonicID is that if one day a user figures that they need more than three trials to log into their device, i.e., the TPR of authentication on that day is lower than 33.3%, they can activate the fine-tuning process in which the system collects 15 seconds of data from this user on that day to further fine-tune their trained individual model. This process can happen while the user is using the device after successfully logging into the system without interrupting their normal usage of the device. Generally, the fine-tuning process will only be carried out once so that limited burden of data collection will be added onto the user while the system guarantees a satisfactory authentication experience in the meantime.
Following this principle of fine-tuning, with the models trained on 1-minute data from Day 1, we fine-tuned the models for 4 participants on Day 2 (P32, P33, P36, and P40) and for 3 participants on Day 3 (P10, P27 and P30). The models were fine-tuned for 2 epochs. For the 4 participants with models fine-tuned on Day 2, the average TPR on Day 2 increases from 16.9% to 74.3% while FPR decreases from 4.6% to 3.6%. At the same time, the average TPR for these 4 participants on Day 3 also increases from 33.7% to 57.7% while FPR decreases from 4.8% to 3.1%. For the 3 participants with models fine-tuned on Day 3, the average TPR on Day 3 increases from 20.0% to 63.3% while FPR decreases from 6.7% to 4.7%. These results showcased that the performance of SonicID improves significantly in terms of both TPR and FPR across several days with only 15 seconds of data collected on a second day. Applying this fine-tuning process on the leave-one-day-out evaluation presented in Sec. 6.3, we obtained the updated evaluation results in Tab. 3. Tab. 3 displays that BAC is improved to around 90% for the across day evaluation, with the fine-tuning process applied on only a few participants, which proved the feasibility of adopting fine-tuning to improve the performance of SonicID across days.
Table 3. Average TPR, FPR and BAC across 40 Participants with Fine-tuning Process Applied on Selected Participants. Data Format: TPR/FPR/BAC. Results with Training and Testing Data from the Same Day are in Bold.
Training Data Source Day 1 Day 2 Day 3
Train on Day 1 97.4%/4.3%/96.6% 86.5%/4.0%/91.3% 80.9%/3.5%/88.7%
Train on Day 2 86.4%/3.8%/91.3% 96.5%/3.6%/96.5% 84.1%/3.7%/90.2%
Train on Day 3 81.3%/3.5%/88.9% 84.5%/3.2%/90.6% 98.2%/3.1%/97.5%
6.5. Impact of Different Body Postures
As discussed in Sec. 5, the last 16 participants evaluated the system in various body postures including talking, standing and lying down. To explore the generalizability of SonicID when users are in different postures, we loaded the individual models trained on data collected when participants were sitting still and directly tested them with data collected when these participants were in different body postures. Note that although the positive samples only came from the data that each participant collected in specific scenarios, the negative samples came from the data all other participants collected in all scenarios, including sitting still. Same as Sec. 6.3, leave-one-day-out evaluation was carried out to alleviate randomness across days. We averaged the results under two categories, i.e. performance on the same day of data collection and performance on a different day, and showcased the results in Tab. 4.
Table 4. Average TPR, FPR and BAC across Last 16 Participants for Different Application Scenarios. The Results are Averaged through Leave-One-Participant-Out Evaluation. Data Format: TPR/FPR/BAC.
Scenarios Same Day Evaluation Across Day Evaluation Across Day Evaluation with Fine-tuning
Sitting Still 97.3%/4.2%/96.5% 72.4%/4.2%/84.1% 80.4%/3.9%/88.3%
Talking 92.8%/4.2%/94.3% 68.3%/4.2%/82.0% 74.2%/3.9%/85.2%
Standing 92.2%/4.2%/94.0% 69.7%/4.2%/82.8% 73.9%/3.9%/85.0%
Lying Down 87.4%/4.2%/91.6% 62.9%/4.2%/79.3% 69.1%/3.9%/82.6%
As shown in Tab. 4, in same day evaluation, TPR remains over 90% for talking and standing scenarios and over 85% for lying down scenarios even when the model is trained with data of sitting still only. BAC maintains over 90% for all scenarios. In across day evaluation, FPR remains consistent under 5% while TPR decreases compared with same day evaluation for all scenarios. BAC maintains around 80% for all scenarios. Note that FPR for all scenarios under one evaluation is the same because the negative samples used to evaluate the model are the same. However, the negative samples used to evaluate the models in different evaluations are different. FPR remaining unchanged indicates that SonicID performs consistently well to reject unknown attackers across different days. With the fine-tuning process in Sec. 6.4 applied, TPR across days improves by around 5% and BAC is increased to around 85% for all scenarios. Note that fine-tuning data only includes data collected when users are sitting still as well. These results proved that SonicID generalizes to various application scenarios well with only 1 minute of training data collected when the users are sitting. Specifically, we think that the selection of static instances contributed to the fact that the speech-related facial movements did not have a severe impact on the system performance because there were still relatively static moments, e.g. pauses between words, phrases and sentences, in each session even when the participants kept talking and the non-static data was not used for authentication in SonicID.
Table 5. Average TPR, FPR and BAC across 40 Participants with Different Channels of Signals. Models are Trained with Data from Day 1. Data Format: TPR/FPR/BAC.
@@ 6.5. Impact of Different Body Postures
Both Sides 97.4%/4.3%/96.6% 80.8%/4.1%/88.4% 75.2%/3.8%/85.7%
Right Side 90.2%/3.6%/93.3% 63.3%/3.6%/79.9% 52.7%/3.6%/74.5%
Left Side 86.0%/5.8%/90.1% 44.0%/5.4%/69.3% 31.1%/4.7%/63.2%
6.6. Ablation Study
6.6.1. Comparison of Two Sides of Sensors
When we designed the prototype for the SonicID system, we instrumented one pair of speaker and microphone on each side of the glasses to collect complete biometric information from users’ face. However, many commercial smart glasses only have the speaker and microphone on one side of the device. As a result, we would like to experiment on how the system performs when only one side of sensors are utilized. As designed in Sec. 3, four speaker-microphone channels are input into the deep learning model when two pairs of speakers and microphones are deployed. If we only want to use one pair of speaker and microphone on either right side or left side of the glasses, we can just input one speaker-microphone channel related to this pair of speaker and microphone into the model. The experiment results are demonstrated in Tab. 5.
@@ 6.6.1. Comparison of Two Sides of Sensors
6.6.2. Impact of Pre-training Stage
Sec. 3.4 introduced the two-stage machine learning pipeline we used to train the models and evaluate SonicID. In order to validate that the pre-training stage actually plays a positive role in the system, we run another experiment of directly training the feature extractor and binary classifier without the pre-training stage. All other parameters were kept the same and the models were trained on 1-minute data from Day 1 for each participant for 30 epochs in total. The average TPR/FPR/BAC for Day 1, Day 2 and Day 3 in this case are 96.6%/4.0%/96.3%, 75.7%/3.7%/86.0% and 68.4%/3.3%/82.6% respectively while those for the original evaluation are 97.4%/4.3%/96.6%, 80.8%/4.1%/88.4% and 75.2%/3.8%/85.7%. The comparison of the results displayed that the pre-training stage results in better BAC and TPR, especially on Day 2 and Day 3. This aligns with our intention of designing the pre-training stage in Sec. 3.4, which is to better extract unique features associated with each participant that can help improve performance across days.
6.6.3. Selection of Static Instances
In the data processing pipeline of SonicID, the 10 most static 0.06-second instances are picked from each 10-second session to train the model and evaluate the system. These parameters were empirically selected in the pilot study. To showcase that the selection leads to reasonable performance of the system, we run experiments on using 5 and 15 most static instances to authenticate participants. Moreover, we experimented on different lengths of instances, including three 600-sample frames, i.e., 0.036 seconds of data and ten 600-sample frames, i.e., 0.12 seconds of data. The experimental results are displayed in Tab. 6.
For all other selections of static instances except the combination we used in our evaluation, both TPR and FPR decreased on Day 2 and Day 3. In other words, the system was more strict in authenticating users. For the selection of 5 instances and 0.036-second instances, we believe that this was because only perfectly static instances were included in training and the model did not see enough variance so it could not be generalized to various scenarios well. For the selection 15 instances and 0.12-second instances, we assume that this was because the model was overfitted to the training data on Day 1 since the TPR on Day 1 was better but it failed to generalize across different days well. Despite of the variation among different selections of static instances, there was only slight divergence in terms of the balanced accuracy. As a result, the selection of ten 0.06-second static instances in our main study evaluation reached a trade-off balancing TPR and FPR.
Table 6. Average TPR, FPR and BAC across 40 Participants with Different Selections of Static Instances. Models are Trained with Data from Day 1. Data Format: TPR/FPR/BAC. The Combination used in Main Study Evaluation is in Bold.
Selection of Static Instances Day 1 Day 2 Day 3
0.06 seconds, 5 Instances 96.7%/3.8%/96.5% 75.7%/3.5%/86.1% 70.9%/3.2%/83.9%
0.06 seconds, 15 Instances 98.0%/4.2%/96.9% 78.6%/4.0%/87.3 71.1%/3.5%/83.8%
0.06 seconds, 10 Instances 97.4%/4.3%/96.6% 80.8%/4.1%/88.4% 75.2%/3.8%/85.7%
0.036 seconds, 10 Instances 98.1%/4.3%/96.9% 79.7%/4.2%/87.8% 70.0%/3.5%/83.2%
0.12 seconds, 10 Instances 97.5%/3.6%/97.0% 78.1%/3.4%/87.3% 72.5%/3.1%/84.7%
@@ 6.7. Usability
A questionnaire was distributed to participants after the user study to collect their demographic information and feedback on the prototyped wearable device. First, the participants evaluated the overall experience of wearing the device by answering the question "How comfortable is this wearable device to wear around the face? (0 most uncomfortable, 5 most comfortable)". Among 40 participants, the average rating for this question is 3.4 (std=0.9), indicating that the device with the SonicID system is overall comfortable to wear around the face. Next we specifically asked participants to evaluate the weight of the device with the question "How acceptable do you find the weight of our wearable device? (0 most unacceptable, 5 most acceptable)". The average rating is 3.8 (std=0.9), validating that the weight of the device does not cause much burden on users.
In the SonicID system, we utilized two ultrasonic signals in the frequency ranges over 18 kHz. Theoretically, most people are not capable of hearing the sounds emitted from the system. However, due to the limitation of the hardware components, there might be some frequency leakage into the lower frequency range and some users might be able to hear the sound. As a result, we asked the participants to also answer the question "Can you hear the sound emitted from our system?". 29 out of 40 participants answered "No" to this question while the other 11 participants reported "Yes". For those participants who answered "Yes" to this question, a follow-up question "If yes, how comfortable do you feel about the emitted sound? (0 most uncomfortable, 5 most comfortable)" was asked. The average rating for this question among 11 participants is 4.4 (std=0.5). Even though some users might hear the sound from the system, they generally find this sound not bothering them a lot. Since the echo profile we use as an instance to authenticate users is as short as 0.06 s as discussed in Sec. 3.3, we do not expect the sound to trouble the users continuously. Furthermore, it is possible to lower the strength of the transmitted signals to make them less noticeable by users.
@@ 7.1. Power Consumption
We wanted to evaluate the power consumption of the SonicID system in operation. For this purpose, we measured the current that flowed through system with a current ranger [^5] https://lowpowerlab.com/guide/currentranger/ when the system was operating. The current was measured at 165.4 mA with a voltage at 3.3 V. This gave us an average power consumption of 545.8 mW. Considering that SonicID only needs an instance of 0.06 s to authenticate users every time, theoretically it consumes 32.7 mJ of energy for each authentication trial. Note that the calculation of the power consumption here does not include that of the machine learning inference because the inference stage is currently carried out on a local server. To implement a real-time pipeline for SonicID, we could run the model inference stage on a low-power chip powered with a CNN accelerator, such as MAX78002 (Devices, 2024). Prior work (Li et al., 2024b) presented that ResNet-18 can be run on MAX78002 to make 30 inferences per second with a power consumption of 79.0 mW. Hence, we anticipate that the machine learning pipeline of SonicID can also be deployed on MAX78002 to realize a real-time pipeline since we also use a model based on the ResNet-18 architecture. Under this circumstance, SonicID needs 0.093 s in time and 58.1 mJ in energy to perform each trial of user identification. If SonicID is deployed on Meta Ray-Ban with a battery capacity of 154 mAh, the authentication system can be activated around 31,500 times in theory, if SonicID is used alone. However, we believe that these metrics need to be measured more accurately when a real-time pipeline is actually implemented. We leave this for future work.
@@ 7.4. Potential Methods to Improve Performance
As evaluated in Sec. 6, although the FPR is consistently below 5% across several days in evaluation, the TPR reduces on Day 2 and Day 3 compared with Day 1 for some participants. We believe this is due to several reasons. Firstly, the positive samples we collected from each participant every day were just from the 5 minutes of data. The small size of the data could lead to fluctuations in the evaluation results in TPR. Secondly, we noticed that TPR is especially low for P30 and P32 who have long hair and bangs. The hair could sometimes cover the system and prevent the participant from putting the glasses back to the same position on their head every time they remounted the device. We believe that the same reason caused P4 and P27 to have a worse FPR than other participants. In future, we plan to explore several potential methods to improve the performance of the system.
@@ 7.4.1. Explore Movement-based Authentication
In Sec. 6, we removed information of blinking and other movements from the data during training and testing to guarantee a valid facial scanning outcome. However, this information can also be taken advantage of to authenticate users. For instance, the blinking information alone can be exploited as a unique biometric pattern because different users blink differently. Furthermore, we can explore other movement-based authentication methods to add another layer of protection. Users of our system can select a specific facial gesture (e.g. smile, tongue out, blink, etc.) as their password to the system. Our system recognizes both their password and their biometric information when performing this password to authenticate users. This kind of two-layer authentication systems make the system more secure and are possible to implement given that there are already some work validating that deploying acoustic sensing on wearables can help track users’ facial movements (Li et al., 2022; Zhang et al., 2023; Xie et al., 2021; Li et al., 2024a). SonicID can potentially be integrated into these existing systems to realize user authentication without adding extra hardware.